A sample of an Iron Oxalato complex salt weighting 0.13 grams requires 31.27 mL of 0.013 M KMnO4 to turn the solution a very light pink color at the quivalence point.

1.Calculate the number of moles of KMnO4 added.

2. Calculate the number of moles of C2O42− in the sample of the oxalate salt.

3.Calculate the number of grams of C2O42− in the sample.

4. Calculate the percent weight of C2O42− in the original Iron Oxalato Complex salt sample.
Answer #1

Given data,

Mass = 0.13 g

Volume = 31.27 mL

Let us consider a reaction,

5 C2O4^2- (aq) +2 MnO4^- (aq) + 16 H^+ (aq) ---> 10 CO2 (g) + 8 H2O (l) + 2 Mn^2+(aq)

From the reaction,

5 mol C2O4^2- = 2 mol MnO4^-

1)

No of mol of KMnO4^- consumed = Molarity x volume

= 0.013 * 31.27

= 0.40 mmol

2)

No of mol of C2O4^2- reacted = 0.40 * 5 / 2

= 1mmol

number of moles of C2O42− = 1 * 10^-3 mol

3)

number of grams of C2O42− in the sample = No.of moles x molecular weight

= (1 * 10^-3) * 88

= 0.088 g

4)

percent weight of C2O42− = 0.088 / 0.13 * 100

= 0.676 x 100

= 67.6 %